Session: June 2008 (session 2)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (scalambra at Sat May 31 16:36:00 2008) Success :

Onsala (Michael Lindqvist at Sun Jun 1 09:44:24 2008) Minor failures :
Problem with "noisy" RCP channel back (see em059c). Have not been able to find any problems with the receiver (yet). In addition, difficult to get correct values on the BBCs due to our bad/large IF-slope :-(.

Effelsberg (Uwe Bach at Mon Jun 2 08:38:21 2008) Minor failures : LATE
Problems with the VLBA VME System. Changed back to MKIV. Should be okay after 11 UT!

Hartebeesthoek (Jonathan Quick at Mon Jun 2 16:10:31 2008) Success :
Ho know problems - perfect weather

Hartebeesthoek (Jonathan Quick at Mon Jun 2 16:10:45 2008) Success :
Ho known problems - perfect weather

Jodrell_Bank (Alastair Gunn at Tue Jun 10 23:45:05 2008) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Tue Jun 10 23:45:39 2008) Success :
No known problems.